1, 2, assume that the individual`s utility function is given by:

Dp8d0wf21Da-B2oSNg05W_GlfeLrWW6jSbK0hcTq

  1. An individual with an income of 1000 is considering the purchase of a risky asset for $250. There is a 40% chance the asset will earn $1000 and a 60% chance it will be worthless.

  1. Will the individual purchase this asset?

  2. What is the cost of risk for this individual if the asset is purchased?

  3. How much will the cost of risk be if a 2nd identical person is brought in to help purchase the asset?

  1. Another individual is considering purchasing insurance to protect his or her home from a potential flood. There is a 20% chance of a flood. If the flood occurs, it will cost $25,000 to repair the damage. The individual begins with $60,000 in income.

  1. If the individual is offered insurance at a price of $5000, will he or she purchase it?

  2. What is the maximum this individual would be willing to pay for insurance?

  3. Repeat part b) for damages of $5000 and $50000. Explain the different prices in the context of a risk averse individual.
Answer #1

Solution: Given U = UI W income = 1000 risky offet = 8250 = 40.3. Ceara $1000) = 60.4 (worthless @ Individual purchase this a© cost of risk be it and identical person is brought into help purchase the affet: It will remain unchanged. [ (2) Giren o=FIIf individual is offered insurance at price of $5000, he will purchase it since the actuarially fair premium is the expectedExpecfed utility = 60.2) (187.08) +0.8) (R44.99 = 37.416 +195-952 • 0 = 933 368 Certainity equivalent is given af 233.368 = (Certainity equivalent is given as: 215.952 = (I) 0.5 :. I = 46635.2663 ?: max amt of insurance will be paying = 60000 4.6635